OpenStudy (anonymous):

Solve the equation: 4x2 = 36 {6, -6} {3, -3} {2/3, -2/3} {3/2, -3/2}

OpenStudy (saifoo.khan):

therefore: {6,-6}

OpenStudy (anonymous):

{-3 ,+3}

OpenStudy (akshay_budhkar):

{3, -3} you can substitute and verify

OpenStudy (saifoo.khan):

4x^2 - 36 = 0 4(x^2 - 9) = 0 x^2-9 = 0 (x-3)(x + 3)

OpenStudy (anonymous):

4x^2=36 x^2=36/4 x^2=9 x=sqrt9 x=-3,3

OpenStudy (saifoo.khan):

therefore: {3,-3}